# Generative AI for Immersive Communication: The Next Frontier in Internet-of-Senses Through 6G

### v1

@@ IV-B Proposed Architecture for GenAI Enabled Immersive Communication
Concurrently with the 3D view, the VR headset and other wearables receive a multimedia description file containing values related to temperature and vibration. These values are derived from the environmental image and UAV movements, estimated using the LLM. The haptic feedback and potential heat dissipation wearable replicate the estimated environment concurrently with the view, minimizing synchronization latency as much as possible. Notably, since the generated code represents an animation, it is not necessary to update the view at a high frequency. However, the virtual camera in the spherical projection moves according to the drone’s position, continuously received from the Message Queuing Telemetry Transport (MQTT) broker by the user.
@@
Unmanned Aerial Vehicle. The UAV functions as the real-world data capture device, employing its mounted 360 camera to provide live feedback in the form of a 360 video to the VR user. Simultaneously, the UAV performs object detection using YOLOv7, trained on the MS COCO Dataset [21], and subsequent captioning using Inception-v3 [22] on one frame every 30 frames. The resulting object annotations and sensorial data, including position, are transmitted to the cloud, specifically to a LLM for additional contextualization and detailed description. Furthermore, the UAV independently streams its position data, comprising altitude, latitude, and longitude, through Mavlink telemetry messages to a drone Application Programming Interface (API). This information is then relayed to the HMD via the MQTT broker. It is important to highlight that we are not hosting an LLM on the UAV due to its substantial size, computing demands, and energy consumption. Doing so would significantly reduce flight times.
@@
Cloud server. The cloud server primarily hosts Hypertext Transfer Protocol (HTTP) APIs connected to two LLMs, specifically the first LLM, GPT-3.5, and the second LLM, GPT-4. Consequently, the first LLM is responsible for providing more context (enhanced image captioning) from an image caption and its object annotation, received from the UAV. Subsequently, the second LLM is prompted with the generated description from the first LLM as instruction and is tasked with generating Hypertext Markup Language (HTML) code using the A-frame framework [23] to produce immersive 3D WebXR content representing the image description in a 3D space.
@@
It is worth noting that we opt not to directly feed captions from the UAV to the second LLM. Instead, we utilize the first LLM to fuse captions, annotations, and UAV sensor data. This results in captions that are more detailed compared to standard ones [24]. Furthermore, by leveraging the prompt history stored in the LLM memory, we further enhance the accuracy of the descriptions with the LLM in-context learning capability. Thereafter, we utilize a second LLM for code generation, ensuring that it does not impact the memory context of the first LLM. This multi-agents approach also enhances response accuracy, with potential improvements exceeding 6% for GPT-3.5.
@@
Edge Cloud. The edge cloud, located in close proximity to the drone, plays a crucial role in three fundamental computations: video streaming and transcoding, message transmission through a publish/subscribe broker, and web serving. In the traditional method of streaming 360 videos, a media server is utilized to receive an Real-Time Streaming Protocol (RTSP) video stream of equirectangular projected frames. Subsequently, these frames are transcoded using an Advanced Video Coding (AVC)/H.264 encoder for re-streaming through Web Real-Time Communication (WebRTC), as illustrated in [25].
TABLE I: Testbed’s parameters and values.
Parameter Value
Wifi (upload/download) 100Mbps/200Mbps
5G (upload/download) 50Mbps/200Mbps
Ethernet connection 900Mbps/800Mbps
Edge Server CPU 8 cores @ 2.5GHz
Edge Server memory 16GB
Onboard computer memory 8GB
Onboard CPU 4 cores @ 1.5GHz
Distance UAV to Server 300m
Distance VR HMD to Edge Server 100m
VR headset Oculus Quest 2
@@
In contrast, in our proposed architecture centered on Generative AI-driven semantic communication, we employ WebXR code generated using the LLM, specifically GPT-4, to represent virtual 3D objects and multimodal descriptions. This data is then transmitted to the user through a MQTT broker and stored in the edge server to construct a Digital Twin (DT) of the environment. An important consideration is that we refrain from hosting the LLMs at the edge due to their large size and computing requirements.
@@ 10 UAV video of Aalto University Finland 2mins
Network and used hardware. Following the global architecture, the testbed comprises a UAV equipped with a 5G modem for communication, an edge HMD with a Wi-Fi connection (chosen due to the operator’s indoor location), an edge cloud connected through fiber, and a cloud server connected via fiber. All components are located in Finland within distances less than 1 km from each other, except for the cloud server situated in the USA. Table I provides a detailed description of the hardware used.
@@
V-B Prompts and output
At first, the first LLM is prompted by annotations and objects from the UAV and generates the following description for the video taken during our experiment: The image depicts a large red building with a flat roof, surrounded by snow-covered trees and a snow-covered ground. There are two people in the foreground, one of them is holding a camera, and the other appears to be flying a drone.
@@ V-B Prompts and output
You are an assistant that teaches me Primitive Element tags for A-Frame version 1.4.0 and later. Create a ’Description from first LLM .
As an output, the rendered code is represented in Fig. 4, which shows both the 3D content generated based on HTML code created by the LLM and the captured 360 frames from the camera are represented. The view undergoes transformation onto a spherical projection to align with the user’s Field of View (FoV).
@@
Figure 5: Comparison of Bandwidth requirements and latencies
V-C Experimental measurements
To measure bandwidth consumption during the upload phase of traditional video streaming, we recorded the bitrate using the FFmpeg [26] command at the UAV. Simultaneously, we utilized the WebRTC statistics API at the HMD level. For our proposed method, we calculated the average size of the description sent from the UAV and the size of the received LLM-generated code at the HMD. In both traditional and our proposed systems, we analyzed various delays, including the E2E traditional video streaming latency (L). This latency is constituted by the RTSP video stream from the UAV to the edge server, the WebRTC stream from the edge server to the HMD, and the frame rendering delays at the HMD, as depicted in Equation (1).
@@ V-C Experimental measurements
The constituting latencies in this latter case have been measured at the edge server, namely for the RTSP streaming, and at the HMD for WebRTC streaming and rendering. In our method, it mainly encompasses the latency of text prompt to 3D WebXR code generation from the two LLMs used, the code transmission through MQTT, and WebXR code rendering. Considering that we can achieve real-time 30 Frames Per Second (FPS) object detection [27] using the onboard computer and that captioning is only applied to one frame out of 30, we consider the object detection latency negligible. The E2E latency ($\text{L}_{\text{Our Method}}$) can be expressed as shown in Equation (2).
@@ V-D Results and analysis
We observe that in traditional video streaming, the uplink and downlink are almost similar, with the downlink being slightly lower due to WebRTC adapting to the available network bandwidth and latency. This difference in bandwidth requirements is attributed to the complexity of frames based on the videos’ contents. A similar variation is present in the uplink annotation streaming from our method, which is also due to the different descriptions and detected objects from the videos’ frames. The downlink, on the other hand, consistently has the same size, attributed to the code and output of the LLM, which is restricted by the prompt to a predetermined size of generated tokens.
@@
V-E Validation
To validate our proposed system, we compared the descriptions generated from a VLM, namely GPT-4, for the virtual 3D generated frames and the equirectangular frames of the 10 videos. Since direct comparison between frames is not feasible using traditional methods such as Peak Signal-to-Noise Ratio (PSNR), we employed average semantic comparison through Bidirectional Encoder Representations from Transformers (BERT), which is widely used to measure the degree of semantic textual similarity, as depicted in the results in Fig. 6. It is worth noting that the maximum BERT similarity score represents the highest probability of 1. Results generally show a good representation of code-based descriptions, with a maximum matching score of 83% observed for the 8th video representing a Buddhist temple and a minimum score of 43% for the 6th video depicting a park with animals and a lake. The lower quality representation in videos 3, 5, and 6 can be attributed to their complexity and rich set of objects. The WebXR framework used, namely A-Frame [28], generates 3D representations using basic geometric objects, making it challenging to recreate such complex scenes. Additionally, the utilized language models (LLMs) are not fine-tuned to be experts in the WebXR framework. It is worth noting that we have also explored the Code LLaMa, designed specifically for coding tasks. However, this latter’s generated code was very weak compared to GPT-4.
12345678910000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.780.780.780.780.680.680.680.680.550.550.550.550.610.610.610.610.440.440.440.440.430.430.430.430.750.750.750.750.830.830.830.830.730.730.730.730.760.760.760.76Video NumberDescription similarity %
@@ V-E Validation
VI Challenges & Open Research Directions
VI-A Multi-user & Scalability
Scalability for such applications as the one proposed in the use case is quite challenging since the response from an LLM when prompted by multiple tasks might be degraded by up to 3% less accuracy for 50 simultaneous prompts [29]. A scalable 6G network is needed in order to accommodate a large number of immersive users, while dynamically being able to adjusting its resources and services to serve varying demands without compromising performance, reliability, or user experience.
VI-B Latency & Real-time Processing
@@
VI-C Edge Computation Limitations
@@ VI-E Integration & Interoperability
VII Conclusion
In this paper, have laid a foundational framework for integrating LLMs with the IoS within the context of 6G. We have defined the principles of IoS and proposed promising use-cases to demonstrate the potentials of LLMs in realizing low-latency immersive media communication. Within the developed use-cases, we have adopted LLMs as compressors, and we have presented a use case with a real testbed leveraging GenAI for IoS. The measurement methods and analysis of the proposed system have been highlighted, and benchmarked with the traditional way of streaming immersive video. The results showed that LLMs are very effective compressors, leading to significant bandwidth savings. However, their response latency is high for real-time applications. To address these latency issues, an approach aiming at fine-tuning the LLM and hosting it near the user has been designed and presented. As a future work, we aim to explore the use of the fine-tuned LLM on the UAV as an alternative to traditional captioning and object detection.

### v2

@@ V-B Proposed Architecture for GenAI Enabled Immersive Communication
Beyond animated 3D objects, the LLM is able to estimate temperature along wind’s speed and direction thereafter generate a Mulsemedia values map to activate the vibrators of a wearable haptic suit such as the Teslasuit  [^8] https://teslasuit.io/products/teslasuit-4/ and its thermal sensors based on the built-in sensors of the UAV’s flight controller, such as altitude, speed, and Inertial Measurement Unit (IMU). These estimations enable precise replications of environmental conditions, including tactile and thermal feedback, synchronized with the generation of the 3D objects. This creates a comprehensive IoS experience, allowing the user to feel the environmental conditions in real-time, enhancing the immersive quality of the virtual environment.
@@
Concurrently with the 3D view, the VR headset and other wearables receive a Mulsemedia description file containing values related to temperature and vibration. These values are derived from the environmental image and UAV movements, estimated using the LLM. The haptic feedback and potential heat dissipation wearable replicate the estimated environment concurrently with the view, minimizing synchronization latency as much as possible. Notably, since the generated code represents an animation, it is not necessary to update the view at a high frequency. However, the virtual camera in the spherical projection moves according to the drone’s position, continuously received from the Message Queuing Telemetry Transport (MQTT) broker by the user.
@@
Unmanned Aerial Vehicle. The UAV functions as the real-world data capture device, employing its mounted 360 camera to provide live feedback in the form of a 360 video to the VR user. Simultaneously, the UAV’s onboard computer performs object detection on the camera’s captured frames using YOLOv7, trained on the MS COCO  [^9] https://cocodataset.org/ Dataset, and subsequent captioning using Inception-v3 and Long Short-Term Memory (LSTM[28] on one frame every 30 frames. The resulting object annotations resulting from the object detection, brief caption from the LSTM, and sensorial data from the UAV’s embedded sensors, including position data, are transmitted to the cloud in a JavaScript Object Notation (JSON) format, specifically to a LLM for additional contextualization and detailed description. Furthermore, in parallel the UAV independently streams its position data, comprising altitude, latitude, and longitude, through Mavlink telemetry messages to a drone Application Programming Interface (API). This information is then relayed to the HMD via the MQTT broker. It is important to highlight that we are not hosting an LLM on the UAV due to its substantial size, computing demands, and energy consumption. Doing so would significantly reduce flight times.
@@
Cloud server. The cloud server primarily hosts Hypertext Transfer Protocol (HTTP) APIs connected to two LLMs, specifically the first LLM, GPT-3.5-Turbo, which has 175 billion parameters, and the second LLM, GPT-4, which has about 1.8 trillion estimated parameters. Consequently, the first LLM is responsible for providing more context (enhanced image captioning) from an image caption and its object annotation, received from the UAV. Subsequently, the second LLM is prompted with the generated description from the first LLM as instruction and is tasked with generating Hypertext Markup Language (HTML) code using the A-frame  [^10] https://aframe.io/ framework to produce immersive 3D WebXR content representing the image description in a 3D space.
@@
It is worth noting that we employ a multi-agent architecture with two distinct LLMs, each assigned a specific task to enhance the accuracy of responses especially that LLMs might not perform well when dealing with longer text sequences or tasks that require long term planning. This strategy avoids directly feeding captions from the UAV to the second LLM. Instead, the first LLM fuses captions, annotations, and UAV sensor data, resulting in more detailed captions compared to standard ones Furthermore, by leveraging the prompt history stored in the LLM memory, we enhance the accuracy of the descriptions through the LLM’s in-context learning capability. Subsequently, we utilize a second LLM for code generation, ensuring that it does not impact the memory context of the first LLM. This multi-agent approach has been shown to improve response accuracy, with potential enhancements exceeding 6% for GPT-3.5.
@@
Edge Cloud. The edge cloud, located in close proximity to the drone, plays a crucial role in three fundamental computations: video streaming and transcoding, message transmission through a publish/subscribe broker, and web serving. In the traditional method of streaming 360 videos, a media server is utilized to receive an RTSP video stream of equirectangular projected frames. Subsequently, these frames are transcoded using an Advanced Video Coding (AVC)/H.264 encoder for re-streaming through WebRTC, as illustrated in [14].
@@
In contrast, in our proposed architecture centered on Generative AI-driven semantic communication, we employ WebXR code generated using the LLM, specifically GPT-4, to represent virtual 3D objects and multimodal descriptions. This data is then transmitted to the user through a MQTT broker and stored in the edge server to construct a DT of the environment. An important consideration is that we refrain from hosting the LLMs at the edge due to their large size and computing requirements.
@@ 10 UAV video of Aalto University Finland 2mins
Network and used hardware. Following the global architecture, the testbed comprises a UAV equipped with a 5G modem for communication, an edge HMD with a Wi-Fi connection (chosen due to the operator’s indoor location), an edge cloud server connected through fiber, and a cloud server connected via fiber. All components are located in Finland within distances less than 1 km from each other, except for the cloud server situated in the USA. Table III provides a detailed description of the hardware used.
TABLE III: Testbed’s parameters and values.
Parameter Value
Wifi (upload/download) 100Mbps/200Mbps
5G (upload/download) 50Mbps/200Mbps
Ethernet connection 900Mbps/800Mbps
Edge Server CPU 8 cores @ 2.5GHz
Edge Server memory 16GB
Onboard computer memory 8GB
Onboard CPU 4 cores @ 1.5GHz
Distance UAV to Server 300m
Distance VR HMD to Edge Server 100m
VR headset Oculus Quest 2
@@
VI-B Prompts and output
At first, the first LLM is prompted by annotations and objects from the UAV and generates the following description for the video taken during our experiment:
The image depicts a large red building with a flat roof, surrounded by snow-covered trees and a snow-covered ground. There are two people in the foreground, one of them is holding a camera, and the other appears to be flying a drone.
@@ VI-B Prompts and output
You are an assistant that teaches me Primitive Element tags for A-Frame version 1.4.0 and later. Create a ’Description from first LLM’.
As an output, the rendered code is represented in Fig. 5, which shows both the captured 360 frames from the camera in Fig. 5 (a) and 3D content generated based on HTML code created by the LLM in Fig . 5 (b). The view undergoes transformation onto a spherical projection to align with the user’s FoV.
@@
Figure 6: Comparison of bandwidth requirements and latency scores.
VI-C Experimental measurements
To measure bandwidth consumption during the upload phase of traditional video streaming, we recorded the bitrate using the FFmpeg  [^12] https://www.ffmpeg.org/ command at the UAV. Simultaneously, we utilized the WebRTC statistics API at the HMD level. For our proposed method, we calculated the average size of the description sent from the UAV and the size of the received LLM-generated code at the HMD. In both traditional and our proposed systems, we analyzed various delays, including the E2E traditional video streaming latency (L). This latency is constituted by the RTSP video stream from the UAV to the edge server, the WebRTC stream from the edge server to the HMD, and the frame rendering delays at the HMD, as depicted in Equation (1).
@@ VI-C Experimental measurements
The constituting latencies in this latter case have been measured at the edge server, namely for the RTSP streaming, and at the HMD for WebRTC streaming and rendering. Our method mainly encompasses the latency of text prompt to 3D WebXR code generation from the two LLMs used, the code transmission through MQTT, and WebXR code rendering. Considering that we can achieve real-time 30 Frames Per Second (FPS) object detection using the onboard computer and that captioning is only applied to one frame out of 30, we consider the object detection latency negligible. The E2E latency ($\text{L}_{\text{Our Method}}$) can be expressed as shown in Equation (2).
@@ VI-D Results and analysis
We observe that in traditional video streaming, the uplink and downlink are almost similar, with the downlink being slightly lower due to WebRTC adapting to the available network bandwidth and latency. This difference in bandwidth requirements is attributed to the complexity of frames based on the content of each video. A similar variation is present in the uplink annotation streaming from our method, which is also due to the different descriptions and detected objects from the videos’ frames. The downlink, on the other hand, consistently has the same size, attributed to the code and output of the LLM, which is restricted by the prompt to a predetermined size of generated tokens.
@@
12345678910000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.830.830.830.830.790.790.790.790.620.620.620.620.60.60.60.60.690.690.690.690.780.780.780.780.670.670.670.670.770.770.770.770.830.830.830.830.890.890.890.890.780.780.780.780.690.690.690.690.560.560.560.560.720.720.720.720.620.620.620.620.730.730.730.730.690.690.690.690.70.70.70.70.820.820.820.820.80.80.80.80.870.870.870.870.860.860.860.860.760.760.760.760.790.790.790.790.730.730.730.730.80.80.80.80.690.690.690.690.790.790.790.790.950.950.950.950.910.910.910.91Video NumberBERT semantic similarity scoreOur captioningViT over GPT-2 captioningGPT-4o captioning
Figure 7: Description similarity between generated captions compared to human captions
12345678910000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.811110.780.780.780.780.680.680.680.680.550.550.550.550.610.610.610.610.440.440.440.440.430.430.430.430.750.750.750.750.830.830.830.830.730.730.730.730.760.760.760.76Video NumberBERT semantic similarity score
@@
VI-E Validation
To validate our proposed system, we evaluated the semantic similarity between the output of the first LLM used in our framework and human-annotated descriptions of the 10 video frames. We employed BERT, a widely recognized method for assessing the degree of semantic textual similarity. Additionally, we compared our results with those obtained from captioning methods based on transformer architectures, namely ViT over GPT-2, and GPT-4o. Fig. 7 illustrates the similarity scores between human-generated captions and those produced by our method, ViT over GPT-2, and GPT-4o for a randomly selected frame from each video. The results demonstrate the effectiveness of our method, achieving an average similarity of 71% with human captions. Notably, our approach particularly excels with the 10th video, which incorporates UAV sensorial data. While the MLLM GPT-4o might offer the optimal solution due to its training on a broader range of data, it necessitates streaming entire frames, incurring substantially higher bandwidth consumption compared to our approach.
Subsequently, we compared the descriptions generated by a VLM, specifically GPT-4, for both the virtual 3D frames (generated by the second LLM) and the equirectangular frames of the ten videos. As direct frame comparison using traditional metrics like Peak Signal-to-Noise Ratio (PSNR) is not suitable, we employed BERT-based average semantic comparison, with the results depicted in Fig. 8. Notably, the maximum BERT similarity score signifies the highest probability of 1.
Overall, the results indicate a satisfactory representation of code-based descriptions. A maximum matching score of 83% was observed for the 8th video (Buddhist temple), while the minimum score of 43% was associated with the 6th video (park with animals and a lake). The lower representation quality in videos 3, 5, and 6 can be attributed to their inherent complexity and the presence of numerous objects. The A-Frame WebXR framework used in this study generates 3D representations using basic geometric shapes, potentially limiting its ability to accurately recreate such intricate scenes. Moreover, the utilized LLMs were not fine-tuned for expertise in the WebXR framework. We also explored Code LLaMA [^13] https://ai.meta.com/blog/code-llama-large-language-model-coding/, designed specifically for coding tasks; however, its generated code was notably weaker compared to that of GPT-4.
VII Challenges & Open Research Directions
VII-A Multi-user & Scalability
Scalability for such applications as the one proposed in the use case is quite challenging since the response from an LLM when prompted by multiple tasks might be degraded by up to 3% less accuracy for 50 simultaneous prompts [29]. A scalable 6G network is needed in order to accommodate a large number of immersive users, while dynamically being able to adjust its resources and services to serve varying demands without compromising performance, reliability, or user experience.
VII-B Latency & Real-time Processing
@@
VII-C Edge Computation Limitations
@@ VII-E Integration & Interoperability
VIII Conclusion
This paper has established a foundational framework for integrating LLMs with the IoS within the context of 6G networks. We have defined the key principles of IoS and presented promising use cases that showcase the potential of LLMs in enabling low-latency, multi-sensory communication experiences. Within these use cases, we have explored the application of LLMs as effective compressors and showcased a practical implementation on a real testbed, leveraging generative AI for the IoS. The measurement methodologies and analysis of the proposed system have been meticulously detailed and benchmarked against traditional approaches to multi-sensory data transmission. Our results demonstrate that LLMs can achieve significant bandwidth savings; however, their response latency currently presents a challenge for real-time applications. To alleviate this limitation, we have designed and presented an approach focused on fine-tuning LLMs and deploying them closer to the user. Looking ahead, we intend to investigate the use of fine-tuned LLMs directly on UAVs as an alternative to conventional captioning and object detection methods, potentially enhancing the sensory experience within IoS applications.